# T-Shirt: Token-Selective Hierarchical Data Selection for Instruction Tuning

### v1

@@ 5 Conclusion
In this paper, we present Token-Selective Hierarchical Data Selection for Instruction Tuning (T-Shirt), a framework for selecting high-quality subsets from instruction tuning datasets. Our method introduces S-IFD, a quality metric that selectively incorporates only informative response tokens into quality assessment, and a hierarchical selection strategy that favors samples surrounded by consistently high-quality neighbors. Extensive experiments across a range of base LLMs and instruction tuning datasets demonstrate that T-Shirt is both effective and scalable, consistently outperforming existing baselines. At the same time, it remains cost-effective and computationally efficient. These results underscore the importance of fine-grained and robust quality evaluation for data selection. We discuss the limitations and future work in Appendix B.
Acknowledgments
We thank Peng Cao, Xianhang Cheng, and Yuehui Qian for their valuable comments on this draft.
@@ Acknowledgments
  • Section B: Limitations and Future Work
  • Section C: Experimental Setup Details
    • Section C.1: Instruction Tuning Datasets
    • Section C.2: Baselines
    • Section C.3: Benchmarks
    • Section C.4: Implementation Details of T-Shirt
    • Section C.5: Instruction Tuning Details
@@ Appendix A Summary of Contributions
  • We analyze existing instruction tuning datasets and find that over 20% of the response tokens are uninformative. Excluding these uninformative tokens from IFD score computation significantly alters the results (see Section 3.1).
  • We investigate the robustness of IFD and observe that it is sensitive to small, semantics-preserving perturbations. Our findings suggest that a high IFD score sometimes reflects superficial features of the training sample rather than its intrinsic quality (see Section 3.2).
  • To address these limitations, we present T-Shirt, a novel data selection framework. We first introduce S-IFD, a new quality metric that incorporates token-level informativeness into assessment. In addition, we propose a robust, hierarchical selection pipeline that prioritizes samples with high average neighborhood quality and low variance (see Section 3).
  • We conduct extensive experiments showing that our method achieves superior performance using no more than 5% of the full instruction tuning datasets. Our approach consistently outperforms prior data selection methods across varying dataset sizes and base LLMs. Moreover, it remains cost-effective and efficient: on a single GPU, it processes a 52k-sample dataset in approximately 40 minutes, with no additional API costs (see Section 4).
Appendix B Limitations and Future Work
While our method shows promising results in data selection for instruction tuning, it has several limitations, which we outline along with directions for future work:
  • Model Scale  Due to computational constraints, our experiments are limited to models in the 7B–8B parameter range as is common in several published papers in instruction tuning [10, 15, 20]. We will apply our data selection method to larger models in future work.
  • Data Security Considerations  Our approach focuses solely on identifying high-quality subsets of instruction tuning datasets and does not address security concerns. Future work could extend our method to incorporate data selection strategies that promote safety and trustworthiness in LLMs.
  • Theoretical Foundations  Our hierarchical selection method is inspired by research in adversarial machine learning [46, 47], robustness measures [48, 49], and LLM robustness [50, 33], and empirical results support its effectiveness. However, a deeper theoretical understanding is still lacking. In future work, we aim to establish a stronger theoretical foundation to better explain its behavior and potentially improve its performance.
Appendix C Experimental Setup Details
C.1 Instruction Tuning Datasets
@@
  • Alpaca-GPT-4 [11] contains 52k instruction-response pairs. It keeps the original instructions from the Alpaca dataset [45], but regenerates the responses using GPT-4 [2] instead of GPT-3.5, resulting in improved response quality.
  • Magpie [10], specifically the Magpie-Pro-300K-Filtered version, [^2] https://huggingface.co/datasets/Magpie-Align/Magpie-Pro-300K-Filtered is a fully synthetic dataset comprising 300k high-quality instruction-response pairs. It is constructed by prompting Llama-3-70B-Instruct to generate one million samples, followed by multi-stage filtering. Compared to Alpaca-GPT-4, Magpie is not only larger but also of much higher quality.
C.2 Baselines
@@
  • Longest [23] applies a simple heuristic that selects the longest samples from the dataset, based on the assumption that longer responses tend to be more informative and useful for instruction tuning. We tokenize responses using the Llama-3.1-8B tokenizer and select those with the highest token counts.
  • Deita [14] follows a four-step selection process. It begins with a small seed dataset and evolves its samples to span different levels of quality and complexity, similar to WizardLM [12]. Then, it prompts ChatGPT to label these samples by quality and complexity. Using the labeled data, it trains its own LLM-based scorers. Finally, it selects high-scoring, diverse samples from the full dataset, skipping those similar to ones already chosen. In our experiments, we use their released scorers and codebase to select data.
  • DS2 [15] also follows a four-step selection process. It first scores the full dataset by prompting GPT-4o-mini. Then, it learns a score-transition matrix to correct possible scoring errors. Next, it estimates each sample’s diversity using feature embeddings by measuring the average cosine similarity to its $K$-nearest neighbors. Finally, it selects samples with both high corrected quality scores and high diversity. In our experiments, we use GPT-4o-mini-2024-07-18 and DS2’s codebase to score and select the data.
C.3 Benchmarks
We extensively evaluate instruction-tuned LLMs using eight benchmarks, including six standardized ones from the OpenLLM Leaderboard [34, 35]:
  • ARC-Challenge  [36]: a challenging subset of the multi-choice question answering benchmark ARC, consisting of 2,590 natural science questions.
  • HellaSwag  [37]: a commonsense inference benchmark with 10k test samples, designed to assess models’ ability to complete sentences plausibly.
  • MMLU  [38]: a diverse question-answering benchmark covering 57 tasks (e.g., STEM, US history), used to measure models’ world knowledge and problem-solving skills.
  • TruthfulQA  [39]: a benchmark of 817 questions focused on assessing models’ truthfulness in the face of common misconceptions and false beliefs.
  • BBH  [40]: a collection of 23 challenging tasks from BIG-Bench [51], often requiring multi-step reasoning.
  • GSM8k  [41]: a test set of roughly 1k grade-school-level math word problems targeting multi-step mathematical reasoning.
Following the OpenLLM Leaderboard, we use the codebase of LM-Evaluation-Harness [42] for evaluation. We report normalized accuracy for ARC-Challenge, HellaSwag, and BBH; accuracy for MMLU and TruthfulQA; and exact match for GSM8k.
@@ C.3 Benchmarks
  • Arena-Hard  [43]: comprises 500 challenging queries sourced from Chatbot Arena [52] and WildChat-1M [53]. It introduces the style-controlled win rate (SC) to mitigate judges’ bias related to stylistic features, including answer length and density of markdown headers. We use GPT-4-0314 as the baseline model and Gemini-2.5-Flash-Preview-04-17 as the judge, reporting SC scores.
  • AlpacaEval 2.0  [44]: contains 805 queries from AlpacaEval [45] and uses the length-controlled win rate (LC) to mitigate judges’ bias on response length. We use GPT-4-1106-Preview as the baseline model and GPT-4o-2024-08-06 as the judge, reporting LC scores.
@@
C.4 Implementation Details of T-Shirt
To implement hierarchical selection in our method, a straightforward approach would be to iterate $M$ times and recompute the S-IFD score for each perturbed sample individually, as shown in Algorithm 1. However, this is computationally inefficient, especially since GPT-2 is relatively small and leaves much of the GPU unused. Instead, we generate a noise tensor of shape $M\times(L+T)\times d$, where $M$ is the number of perturbations, $L$ is the instruction length, $T$ is the response length, and $d$ is the embedding dimension. This noise is added to the sample’s token embeddings using broadcasting. This allows us to compute perplexities for all $M$ perturbed samples in a single forward pass through GPT-2, which significantly reduces the time cost compared to repeated evaluations with a for-loop.
C.5 Instruction Tuning Details
For data selected from Alpaca-GPT-4, we follow the training settings used in prior works [16, 17, 45]: a learning rate of 2e-5, 3 training epochs, a batch size of 128, a warmup ratio of 0.03, and a cosine learning rate schedule. For data selected from Magpie, we keep the same settings but reduce the number of training epochs to 2, in line with the official configuration of Magpie [10]. All instruction tuning experiments are conducted on a server equipped with four NVIDIA A6000 GPUs.
@@ Appendix D Ablation Study Details
D.1 Ablation on Neighbor Variance in Hierarchical Selection
Figure 5: Ablation study on the impact of preferring samples with low-variance ($\hat{\sigma}^{2}(x,y)$ in Equation 4) versus high-variance neighbors during hierarchical selection.  The results indicate that selecting samples whose neighbors exhibit high variance significantly degrades performance.
We examine the role of selecting samples whose neighbors exhibit low quality variance in the hierarchical selection process (see Algorithm 1 and Section 3.2 for details). As shown in Figure 5, performance for both Llama-3.1-8B and Qwen-2.5-7B drops significantly when high-variance neighborhoods are preferred over low-variance ones. This highlights the importance of selecting samples from consistently high-quality neighborhoods and further validates the effectiveness of our hierarchical selection approach.
D.2 Ablation on Token Selection Ratio
@@
We ablate the token selection ratio $k\%$ to assess its impact on data selection performance. Here, a ratio of 100% corresponds to using IFD (i.e., all tokens are used for computing quality scores), rather than S-IFD. As shown in Figure 6(a), Llama-3.1-8B achieves optimal performance at $k=75$, while Qwen-2.5-7B’s performance peaks at $k=50$. These results support our observation in Section 3.1 that over 20% of response tokens in instruction tuning datasets are uninformative. Using S-IFD to focus only on informative tokens enhances data quality assessment and improves data selection for instruction tuning.
D.3 Ablation on Number of Perturbations

### v2

@@ 6 Conclusion
In this paper, we present Token-Selective Hierarchical Data Selection for Instruction Tuning (T-Shirt), a framework for selecting high-quality subsets from instruction tuning datasets. Our method introduces S-IFD, a quality metric that selectively incorporates only informative response tokens into quality assessment, and a hierarchical selection strategy that favors samples surrounded by consistently high-quality neighbors. Extensive experiments across a range of base LLMs and instruction tuning datasets demonstrate that T-Shirt is both effective and scalable, consistently outperforming existing baselines. At the same time, it remains cost-effective and computationally efficient. These results underscore the importance of fine-grained and robust quality evaluation for data selection.
Acknowledgments and Disclosure of Funding
We thank Peng Cao, Xianhang Cheng, and Yuehui Qian for their valuable comments on this paper. This work was partially supported by NSF CAREER 2340006. F. Hamman was also supported by Ann G. Wylie Dissertation Fellowship. Y. Fu gratefully acknowledges support from the NeurIPS 2025 Travel Award.
@@
  • Section B: Experimental Setup Details
    • Section B.1: Instruction Tuning Datasets
    • Section B.2: Baselines
    • Section B.3: Benchmarks
    • Section B.4: Implementation Details of T-Shirt
    • Section B.5: Instruction Tuning Details
  • Section C: Additional Experiments
    • Section C.1: CDF of $|\Delta_{t}|$ with GPT-2 as $\theta^{\prime}$
    • Section C.2: Performance of T-Shirt on Qwen-2.5-14B
    • Section C.3: Standard Deviation of Results
@@ Appendix A Summary of Contributions
  • We analyze existing instruction tuning datasets and find that over 20% of the response tokens are uninformative. Excluding these uninformative tokens from IFD score computation significantly alters the results (see Section˜3.1).
  • We investigate the robustness of IFD and observe that it is sensitive to small, semantics-preserving perturbations. Our findings suggest that a high IFD score sometimes reflects superficial features of the training sample rather than its intrinsic quality (see Section˜3.2).
  • To address these limitations, we present T-Shirt, a novel data selection framework. We first introduce S-IFD, a new quality metric that incorporates token-level informativeness into assessment. In addition, we propose a robust, hierarchical selection pipeline that prioritizes samples with high average neighborhood quality and low variance (see Section˜3).
  • We conduct extensive experiments showing that our method achieves superior performance using no more than 5% of the full instruction tuning datasets. Our approach consistently outperforms prior data selection methods across varying dataset sizes and base LLMs. Moreover, it remains cost-effective and efficient: on a single GPU, it processes a 52k-sample dataset in approximately 40 minutes, with no additional API costs (see Section˜4).
Appendix B Experimental Setup Details
B.1 Instruction Tuning Datasets
@@
  • Alpaca-GPT-4 [peng2023instructiontuninggpt4] contains 52k instruction-response pairs. It keeps the original instructions from the Alpaca dataset [Taori2023alpaca], but regenerates the responses using GPT-4 [openai2024gpt4technicalreport] instead of GPT-3.5, resulting in improved response quality.
  • Magpie [xu2025magpie], specifically the Magpie-Pro-300K-Filtered version, [^2] https://huggingface.co/datasets/Magpie-Align/Magpie-Pro-300K-Filtered is a fully synthetic dataset comprising 300k high-quality instruction-response pairs. It is constructed by prompting Llama-3-70B-Instruct to generate one million samples, followed by multi-stage filtering. Compared to Alpaca-GPT-4, Magpie is not only larger but also of much higher quality.
B.2 Baselines
@@
  • Longest [zhao2024long] applies a simple heuristic that selects the longest samples from the dataset, based on the assumption that longer responses tend to be more informative and useful for instruction tuning. We tokenize responses using the Llama-3.1-8B tokenizer and select those with the highest token counts.
  • Deita [liu2024deita] follows a four-step selection process. It begins with a small seed dataset and evolves its samples to span different levels of quality and complexity, similar to WizardLM [xu2024wizardlm]. Then, it prompts ChatGPT to label these samples by quality and complexity. Using the labeled data, it trains its own LLM-based scorers. Finally, it selects high-scoring, diverse samples from the full dataset, skipping those similar to ones already chosen. In our experiments, we use their released scorers and codebase to select data.
  • DS2 [pang2025ds2] also follows a four-step selection process. It first scores the full dataset by prompting GPT-4o-mini. Then, it learns a score-transition matrix to correct possible scoring errors. Next, it estimates each sample’s diversity using feature embeddings by measuring the average cosine similarity to its $K$-nearest neighbors. Finally, it selects samples with both high corrected quality scores and high diversity. In our experiments, we use GPT-4o-mini-2024-07-18 and DS2’s codebase to score and select the data.
B.3 Benchmarks
We extensively evaluate instruction-tuned LLMs using eight benchmarks, including six standardized ones from the OpenLLM Leaderboard [open-llm-leaderboard-v1, open-llm-leaderboard-v2]:
  • ARC-Challenge  [allenai:arc]: a challenging subset of the multi-choice question answering benchmark ARC, consisting of 2,590 natural science questions.
  • HellaSwag  [zellers2019hellaswag]: a commonsense inference benchmark with 10k test samples, designed to assess models’ ability to complete sentences plausibly.
  • MMLU  [hendrycks2021mmlu]: a diverse question-answering benchmark covering 57 tasks (e.g., STEM, US history), used to measure models’ world knowledge and problem-solving skills.
  • TruthfulQA  [lin2022truthfulqa]: a benchmark of 817 questions focused on assessing models’ truthfulness in the face of common misconceptions and false beliefs.
  • BBH  [suzgun2022bbh]: a collection of 23 challenging tasks from BIG-Bench [srivastava2023beyond], often requiring multi-step reasoning.
  • GSM8k  [cobbe2021gsm8k]: a test set of roughly 1k grade-school-level math word problems targeting multi-step mathematical reasoning.
Following the OpenLLM Leaderboard, we use the codebase of LM-Evaluation-Harness [eval-harness] for evaluation. We report normalized accuracy for ARC-Challenge, HellaSwag, and BBH; accuracy for MMLU and TruthfulQA; and exact match for GSM8k.
@@ B.3 Benchmarks
  • Arena-Hard  [li2024arenahard]: comprises 500 challenging queries sourced from Chatbot Arena [chatbotarena] and WildChat-1M [zhao2024wildchat]. It introduces the style-controlled win rate (SC) to mitigate judges’ bias related to stylistic features, including answer length and density of markdown headers. We use GPT-4-0314 as the baseline model and Gemini-2.5-Flash-Preview-04-17 as the judge, reporting SC scores.
  • AlpacaEval 2.0  [dubois2024alpacaeval2]: contains 805 queries from AlpacaEval [Taori2023alpaca] and uses the length-controlled win rate (LC) to mitigate judges’ bias on response length. We use GPT-4-1106-Preview as the baseline model and GPT-4o-2024-08-06 as the judge, reporting LC scores.
@@
B.4 Implementation Details of T-Shirt
To implement hierarchical selection in our method, a straightforward approach would be to iterate $M$ times and recompute the S-IFD score for each perturbed sample individually, as shown in Algorithm˜1. However, this is computationally inefficient, especially since GPT-2 is relatively small and leaves much of the GPU unused. Instead, we generate a noise tensor of shape $M\times(L+T)\times d$, where $M$ is the number of perturbations, $L$ is the instruction length, $T$ is the response length, and $d$ is the embedding dimension. This noise is added to the sample’s token embeddings using broadcasting. This allows us to compute perplexities for all $M$ perturbed samples in a single forward pass through GPT-2, which significantly reduces the time cost compared to repeated evaluations with a for-loop.
B.5 Instruction Tuning Details
For data selected from Alpaca-GPT-4, we follow the training settings used in prior works [li2024ifd, li2024superfiltering, Taori2023alpaca]: a learning rate of 2e-5, 3 training epochs, a batch size of 128, a warmup ratio of 0.03, and a cosine learning rate schedule. For data selected from Magpie, we keep the same settings but reduce the number of training epochs to 2, in line with the official configuration of Magpie [xu2025magpie]. All instruction tuning experiments are conducted on a server equipped with NVIDIA A6000 GPUs.
Appendix C Additional Experiments
C.1 CDF of $|\Delta_{t}|$ with GPT-2 as $\theta^{\prime}$
Figure 5: CDF of $|\Delta_{t}|$ (Equation˜2) for response tokens in Alpaca-GPT-4, computed using GPT-2.  The red dash line indicates $|\Delta_{t}|=0.01$.
Figure˜5 shows the cumulative distribution function (CDF) of $|\Delta_{t}|$ (as defined in Equation˜2) for response tokens in the Alpaca-GPT-4 dataset, computed using GPT-2 as $\theta^{\prime}$. Notably, 20% of tokens have $|\Delta_{t}|\leq 0.009$, and 50% have $|\Delta_{t}|\leq 0.050$, indicating that a substantial portion of tokens are relatively uninformative even for GPT-2.
C.2 Performance of T-Shirt on Qwen-2.5-14B
Table 5: Performance comparison between our method, T-Shirt, and other baseline methods on Qwen-2.5-14B and Magpie.  We use each method to select 10k samples from Magpie to instruction-tune Qwen-2.5-14B. We use abbreviated benchmark names as shown in Table˜1. The best results are shown in bold, and the second-best are underlined.
OpenLLM Leaderboard LLM-as-a-Judge $\mu_{\textsc{all}}$
ARC-C HS MMLU TQA BBH GSM $\mu_{\textsc{open}}$ AH [^3] While conducting this experiment, Gemini-2.5-Flash-Preview-04-17 became unavailable, so we used Gemini-2.5-Flash instead. AE-2 $\mu_{\textsc{llm}}$
Qwen-2.5-14B
Longest 68.17 84.01 79.23 58.69 61.00 85.22 72.72 34.40 35.01 34.71 63.22
IFD 68.94 84.05 79.07 57.87 61.78 83.02 72.46 33.10 37.15 35.13 63.12
T-Shirt ($k=50$) 68.60 83.90 79.26 58.60 62.04 84.15 72.76 35.30 38.45 36.88 63.79
In the main experiments (Section˜4), we primarily evaluate T-Shirt on models with 7B–8B parameters. Here, we extend the evaluation to Qwen-2.5-14B and compare T-Shirt against Longest and IFD using experimental settings similar to those in Table˜2. The results are shown in Table˜5. On Magpie, T-Shirt consistently outperforms both baselines across $\mu_{\textsc{open}}$, $\mu_{\textsc{llm}}$, and $\mu_{\textsc{all}}$. Notably, it achieves an average improvement of at least 1.75 points on the LLM-as-a-Judge benchmarks.
C.3 Standard Deviation of Results
Table 6: Average performance and standard deviation of Random and T-Shirt across three random seeds on Qwen-2.5-7B and Alpaca-GPT-4.  We use abbreviated benchmark names as shown in Table˜1.
OpenLLM Leaderboard
ARC-C HS MMLU TQA BBH GSM $\mu_{\textsc{open}}$
Qwen-2.5-7B
[…]
Both the baseline Random and our proposed method T-Shirt involve stochasticity. Within our limited budget, we run each method on Qwen-2.5-7B and Alpaca-GPT-4 using three different random seeds, and report the mean and standard deviation of their performance on the OpenLLM leaderboard benchmarks. The results, shown in Table˜6, are consistent with the observations reported in the main paper.
@@ Appendix D Ablation Study Details
Figure 6: Ablation study on the impact of preferring samples with low-variance ($\hat{\sigma}^{2}(x,y)$ in Equation˜4) versus high-variance neighbors during hierarchical selection.  The results indicate that selecting samples whose neighbors exhibit high variance significantly degrades performance.
@@
Table 7: Ablation study on token weights ($w_{t}$ in Equation˜3) for informative vs. uninformative tokens.  We use abbreviated benchmark names as defined in Table˜1. The best results are shown in bold, and the second-best are underlined. The results indicate that completely removing uninformative tokens from the S-IFD computation is more effective than merely upweighting informative tokens.
$w_{t}$ Ratio OpenLLM Leaderboard
ARC-C HS MMLU TQA BBH GSM $\mu_{\textsc{open}}$
Qwen-2.5-7B
$\text{1.5}\!:\!\text{1}$ 65.44 82.37 74.09 61.44 53.83 86.05 70.54
$\text{2}\!:\!\text{1}$ 65.53 82.17 74.19 60.73 53.95 86.81 70.56
$\text{1}\!:\!\text{0}$ (Ours) 66.21 82.39 74.23 61.58 54.21 86.81 70.91
D.1 Ablation on Neighbor Variance in Hierarchical Selection
We examine the role of selecting samples whose neighbors exhibit low quality variance in the hierarchical selection process (see Algorithm˜1 and Section˜3.2 for details). As shown in Figure˜6, performance for both Llama-3.1-8B and Qwen-2.5-7B drops significantly when high-variance neighborhoods are preferred over low-variance ones. This highlights the importance of selecting samples from consistently high-quality neighborhoods and further validates the effectiveness of our hierarchical selection approach.
D.2 Ablation on Token Selection Ratio
We ablate the token selection ratio $k\%$ to assess its impact on data selection performance. Here, a ratio of 100% corresponds to using IFD (i.e., all tokens are used for computing quality scores), rather than S-IFD. As shown in Figure˜7(a), Llama-3.1-8B achieves optimal performance at $k=75$, while Qwen-2.5-7B’s performance peaks at $k=50$. These results support our observation in Section˜3.1 that over 20% of response tokens in instruction tuning datasets are uninformative. Using S-IFD to focus only on informative tokens enhances data quality assessment and improves data selection for instruction tuning.
D.3 Ablation on Token Weights
We ablate the choice of token weights $w_{t}$ used in computing the S-IFD score in Equation˜3. In addition to removing uninformative tokens by setting their weights to 0, we also experiment with upweighting informative tokens by setting the weight ratios between informative and uninformative tokens to 1.5:1 and 2:1. As shown in Table˜7, removing uninformative tokens entirely yields better performance than merely upweighting informative ones.
D.4 Ablation on Number of Perturbations